OpenStudy (anonymous):

Geometry!! Please wait as I finish drawing the problem. a. andgle A b. angle B c. angle C

OpenStudy (anonymous):



OpenStudy (dumbcow):

use Law of Cosines to find angles

OpenStudy (anonymous):

are you familiar with law of cosines ?

OpenStudy (anonymous):

yea such as soh cah toa?

OpenStudy (dumbcow):

not quite \[c^{2} = a^{2}+b^{2}-2ab \cos C\]

OpenStudy (anonymous):

I don't believe I'm familiar with that :/

OpenStudy (dumbcow):

thats ok, its a formula (you can look it up in your textbook) its like the pythagorean thm but for all triangles not just right triangles anyway using this formula you can solve for an angle given all 3 sides

OpenStudy (dumbcow):

\[\cos C = \frac{a^{2} +b^{2} -c^{2}}{2ab}\]

OpenStudy (anonymous):

like sine, cosine, and tangent?

OpenStudy (dumbcow):

not sure what you mean sin/cos/tan are trig ratios related to a right triangle you are going to use the cos ratio in this formula to solve for one of the angles

OpenStudy (anonymous):

Alrighty, but how do I know which number goes under which letter in the formula?

OpenStudy (dumbcow):

ahh well side "a" is across from angle A and so on...

OpenStudy (dumbcow):